how do you solve this equation?
Log base 2x =3 -log base 2 2x

Answer by stanbon(57347)

how do you solve this equation?
Log base 2x =3 -log base 2 2x
----
log2(x) = 3 - log2(2x)
---
log2(x) + log2(2x) = 3
---
log2[x(2x)] = 3
---
log2[2x^2] = 3
2x^2 = 2^3
x^2 = 4
Positive solution:
x = 2
============================
Cheers,
Stan H.
============================